### v1

# Janus: Resilient and Adaptive Data Transmission for Enabling Timely and Efficient Cross-Facility Scientific Workflows

Abstract.
In modern science, the growing complexity of large-scale projects has increased reliance on cross-facility workflows, where institutions share resources and expertise to accelerate discovery. These workflows often involve transferring massive data over wide-area networks. While high-speed networks like ESnet and data transfer services like Globus have improved data mobility, challenges remain. Large data volumes can strain bandwidth, TCP suffers from retransmissions due to packet loss, and traditional fault-tolerance methods like erasure coding introduce significant overhead.
This paper presents Janus, a resilient and adaptive data transmission approach for cross-facility scientific workflows. Janus uses UDP, integrates erasure coding for fault tolerance, and applies error-bounded lossy compression to reduce overhead. This design enables users to balance transmission time and accuracy based on specific needs. Janus also adapts coding parameters to real-time network conditions and uses optimization models to determine ideal configurations. Experiments show that Janus significantly improves data transfer efficiency while preserving fidelity.
Do, Not, Us, This, Code, Put, the, Correct, Terms, for, Your, Paper
ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
As scientific discovery increasingly relies on advanced technologies and instruments, vast amounts of data are being generated at an unprecedented rate. For example, the International Thermonuclear Experimental Reactor (ITER) (ITER, 2022), a project designed to demonstrate fusion energy as a large-scale, carbon-free energy source, is expected to produce 2 petabytes of data daily starting in 2035. Similarly, large-scale simulations such as the Nyx cosmology code (NYX simulation, 2019), used to simulate the evolution of matter in the universe, can generate several terabytes of data every hour on the fastest U.S. supercomputers. While much of this data is analyzed locally, it must often be transferred across wide-area networks (WANs) to facilitate global collaborations. This is particularly true for ITER, which, despite being located in France, enables scientists from around the world to conduct fusion experiments remotely. Afterward, the resulting data needs to be sent to distant researchers for further analysis, sometimes across continents. In cases requiring significant computational power, the data may need to be migrated to HPC systems, such as the U.S. Department of Energy’s supercomputers, to accelerate processing. These cross-facility workflows are not only vital for solving complex scientific problems but also contribute to the movement toward data democratization—the principle that scientific data should be accessible to a wide array of researchers, regardless of their geographic location, institutional affiliation, or available resources, fostering innovation through diverse perspectives. This underscores the critical importance of resilient, efficient, and timely data transmissions.
Despite advancements in high-speed computer network infrastructures like the Energy Sciences Network (ESnet) (ESnet, 2025) and data transmission services such as Globus (Globus, 2025), several challenges remain. First, the volume of scientific data generated by experiments or simulations is immense, and transferring such large datasets across WANs can be time-consuming due to bandwidth constraints, slowing down the pace of discovery. Second, widely used transport protocols like TCP, which rely on retransmission mechanisms to handle packet loss, can significantly reduce transfer efficiency, particularly over long-distance, high-latency networks. As data sizes and transfer distances grow, retransmissions increase delays, undermining the real-time requirements of scientific workflows. Third, while techniques like erasure coding can mitigate retransmissions by offering fault tolerance, they introduce additional overhead during transmission. Furthermore, fixed configurations for fault tolerance are ill-suited to dynamic network environments, where factors like packet loss fluctuate. This static approach can lead to suboptimal performance, resulting in further inefficiencies.
In this paper, we propose Janus, a resilient and adaptive data transmission approach designed to enable timely and efficient cross-facility scientific workflows. To accelerate data transmission over WANs, Janus adopts the UDP protocol instead of TCP. Since UDP lacks a built-in retransmission mechanism to recover from packet loss, Janus incorporates erasure coding to provide fault tolerance. To minimize the overhead of erasure coding, Janus further combines it with error-bounded lossy compression, a technique already in use by scientists to reduce data size without exceeding acceptable error bounds (Ainsworth et al., 2019a; Liang et al., 2018). This approach allows Janus to balance data transmission time and accuracy. Specifically, Janus can address two types of user requirements by solving different optimization problems: 1) If the user requires data with a guaranteed error bound, Janus minimizes data transmission time. 2) If the user needs data within a specified time frame, Janus minimizes the error in the received data. Additionally, as network conditions such as packet loss rate can fluctuate during transmission, applying a static level of redundancy to the data cannot achieve optimal transfer efficiency. Therefore, Janus employs adaptive algorithms that dynamically adjust erasure coding parameters based on real-time network conditions, ensuring more efficient data transmission.
The contributions of this paper are as follows: 1) We develop two optimization models that determine the optimal erasure coding configurations to meet user-defined requirements for both data transmission time and accuracy. These models provide a systematic approach to balancing reliability and efficiency. 2) We design and implement two adaptive data transfer protocols capable of adjusting erasure coding parameters dynamically based on real-time variations in packet loss rates. These protocols ensure optimal performance by minimizing both transmission time and data loss. 3) We present extensive evaluations through simulations and real-world experiments, which demonstrate the effectiveness of our approach in improving the reliability and efficiency of scientific data transfers.
2. Background
2.1. Fault Tolerant Data Transmission
Fault-tolerant data transmission across wide-area networks ensures reliable communication despite challenges like packet loss, corruption, or congestion. TCP, a widely used transport protocol, employs retransmission mechanisms for reliability. It tracks packets with sequence numbers and acknowledgments; if an acknowledgment is not received within a timeout, TCP retransmits the missing packet. However, retransmissions can cause significant delays and variability (Kumar, 1998; Lakshman et al., 2000; Barakat et al., 2000), making TCP unsuitable for low-latency or time-sensitive scenarios.
Checksums (Stone et al., 1998; Stone and Partridge, 2000; Plummer, 1989) are another key component of fault-tolerant systems, used to detect errors in transmitted data. A checksum is computed and sent with the data packet; the receiver recalculates it to check for mismatches. If an error is found, retransmission is initiated if supported. While effective for error detection, checksums do not correct errors and rely on additional mechanisms like retransmissions for recovery.
Erasure coding (Rizzo, 1997; Srouji et al., 2011; Leong and Ho, 2012) offers an alternative approach to fault tolerance in data transmission by proactively correcting errors. It divides data into fragments and adds parity fragments. Even if some fragments are lost, the original data can be reconstructed from the remaining fragments, reducing reliance on retransmissions and improving performance in high-loss networks. However, it introduces computational overhead for encoding and decoding and increases data volume due to the added parity fragments.
2.2. Scientific Data Refactoring
Unlike traditional lossy compressors (Ballester-Ripoll et al., 2019; Chen et al., 2014; Lakshminarasimhan et al., 2013; Lindstrom and Isenburg, 2006; Sasaki et al., 2015; Lindstrom, 2014; Ainsworth et al., 2018, 2019b, 2019a; Liang et al., 2018), which typically adhere to a single user-defined error bound, data refactoring decomposes data into multiple components, allowing flexible reconstruction at different approximation levels. For instance, JPEG2000 (Rabbani, 2002) uses block-based wavelet transformations to support progressive data reconstruction. Similarly, ZFP (Lindstrom, 2014) provides progressive reconstruction but is limited to a fixed-rate mode. Progressive reconstruction has also been studied for cost-accuracy trade-offs in scientific visualizations (Kumar et al., 2014; Pascucci and Frank, 2001; Laney and Pascucci, 2004; Paoluzzi et al., 2004; Shamir and Pascucci, 2001), though many approaches lack guarantees on error bounds, which are critical for preserving scientific features.
pMGARD (Liang et al., 2021) extends MGARD’s multilevel decomposition approach (Ainsworth et al., 2018, 2019b, 2019a) to offer an error-controlled data refactoring framework for scientific data. It leverages multigrid-based decomposition and $L^{2}$ projection to break data into multiple levels, each represented by multilevel coefficients used for reconstruction. The reconstruction error is mathematically bounded, ensuring precise error control for each coefficient. pMGARD employs bitplane encoding for multilevel coefficients, optimizing them into a hierarchical structure for progressive reconstruction. This hierarchical approach allows for efficient reconstruction with error bounds determined by the number of levels used, making pMGARD highly suitable for scientific applications.
3. Optimization in Scientific Data Transmission
This section provides an overview of Janus and introduces two optimization models that enable it to determine the optimal redundancy for data transmission. All mathematical symbols used in the models are listed in Table 1.
Symbol Description
$t$ Latency of transmitting one fragment from sender to receiver
$r$ Number of fragments that can be transmitted per second
$\lambda$ Packet loss rate (number of packet losses per second)
$n$ Number of fragments in each fault-tolerant group
$L$ Total number of levels after refactoring
$l$ Total number of levels sent by the sender
$k_{i}$ Number of data fragments in each fault-tolerant group of level $i$
$m_{i}$ Number of parity fragments in each fault-tolerant group of level $i$
$s$ Size of each fragment
$S_{i}$ Size of each level after data refactoring
$\varepsilon_{i}$ Relative L-infinity error when using level $1,2,\ldots,i$
of the refactored data to reconstruct the original data
$N_{i}$ Number of fault-tolerant groups of level $i$
Table 1. Mathematical symbols used in our models
3.1. An Overview of Janus
To enable efficient transmission of large scientific datasets over wide-area networks, Janus utilizes the low latency of the UDP protocol for rapid data transfer and addresses packet loss with erasure coding and error-bounded lossy compression, allowing recovery without costly retransmissions.
To illustrate how Janus works, consider the example in Fig. 1. As shown in Fig. 1(a), on the sender side, Janus uses pMGARD to refactor the original data into a hierarchical structure with four levels. The sizes of these four levels are denoted by $S_{1}$, $S_{2}$, $S_{3}$, and $S_{4}$, where $S_{1}<S_{2}<S_{3}<S_{4}$. Adding an additional level from top to bottom improves the reconstruction accuracy, with the error bounds decreasing as follows: $\varepsilon_{1}$ for level 1, $\varepsilon_{2}$ for levels 1 and 2, $\varepsilon_{3}$ for levels 1–3, and so on, where $\varepsilon_{1}>\varepsilon_{2}>\ldots>\varepsilon_{L}$. The relative L-infinity error, defined as
(1) $\varepsilon=\frac{\max(\|d-\tilde{d}\|)}{\max(\|d\|)},$
measures the error between the original data $d$ and the reconstructed data $\tilde{d}$, where $\max(|d-\tilde{d}|)$ is the maximum absolute difference and $\max(|d|)$ is the maximum value in the original data. In the worst-case scenario, when no levels are available for reconstruction, the relative error is 1.
Each level is divided into fragments, with erasure coding applied: for every $k$ data fragments, $m$ parity fragments are generated. The total number of fragments, $n=k+m$, remains constant across levels. For example, $k_{1}=4$, $m_{1}=4$; $k_{2}=5$, $m_{2}=3$; $k_{3}=6$, $m_{3}=2$; and $k_{4}=7$, $m_{4}=1$. These data and parity fragments form fault-tolerant groups (FTGs) for each level. All fragments are transmitted via UDP, with each fragment encapsulated in its own UDP packet.
(a) Sender: The original data is refactored into a hierarchical representation, with each level further divided into fixed-size data fragments. Erasure coding is applied to every $k$ data fragments to generate $m$ parity fragments, forming fault-tolerant groups of $k+m$ fragments. All fragments are transmitted via UDP, with each fragment encapsulated in a separate UDP packet.
(b) Receiver: UDP packets may be lost during transmission. In each fault-tolerant group, as long as no more than $m$ fragments are lost, the missing data fragments can still be recovered. However, if more than $m$ fragments are lost, at least one data fragment becomes unrecoverable, leading to corruption of the entire level. In such cases, an approximation of the original data is reconstructed using the remaining hierarchical structure.
Figure 1. An illustration of our approach
During transmission, some packets may be lost, but the parity fragments allow the receiver to recover many lost data fragments without retransmission. As illustrated in Fig. 1(b), for the first level, as long as no more than 4 fragments are lost within each FTG, the data can be fully recovered. Similarly, lost fragments at levels 2 and 3 can be recovered using parity fragments. However, at level 4, if 2 fragments are lost from an FTG, recovery is not possible due to the single parity fragment. If retransmission is not enabled, only the top three levels can be used for reconstruction within error bound $\varepsilon_{3}$. The number of parity fragments at each level can be adjusted based on user requirements.
3.2. Optimization Models
Janus is designed to address two types of user requirements by solving the following two optimization models.
3.2.1. Minimizing Data Transmission Time with a Guaranteed Error Bound on the Received Data
Scientists often require received data, once decompressed, to maintain an error within a specified bound to ensure that compression-induced errors don’t compromise subsequent analyses. If the user-specified error bound is $\varepsilon$ and $\varepsilon_{l}\leq\varepsilon<\varepsilon_{l-1}$, Janus must ensure the transfer of at least the first $l$ levels of the refactored data. When packet loss occurs, lost fragments can be recovered as long as the number of lost fragments does not exceed the parity fragments within each FTG. If packet loss prevents the recovery of enough levels, Janus retransmits the FTGs with unrecoverable losses using passive retransmission, as opposed to TCP’s active retransmission. This approach may require multiple retransmission rounds due to potential additional packet losses during retransmission.
We develop an optimization model for Janus to minimize overall transmission time. The total size of the first $l$ levels is $S=S_{1}+S_{2}+\dots+S_{l}$, and the number of FTGs is $N=\frac{S}{(n-m)s}$. The time for initial transmission is $t+\frac{nN-1}{r}$, and the expected retransmission time depends on the number of FTGs needing retransmission in each round. When the packet loss rate is relatively low, unrecoverable data losses across FTGs can be modeled as independent events. Under this assumption, the number of FTGs with unrecoverable losses in a given round follows a binomial distribution. For instance, if the number of FTGs transmitted in the initial round of transmission is $N$, and the probability of a single FTG experiencing unrecoverable data loss is $p$, the probability that exactly $K$ out of $N$ FTGs experience such losses is given by $\binom{N}{K}p^{K}(1-p)^{N-K}$. The expected number of FTGs that will need retransmission in the first round of retransmission is $\sum\limits_{K=0}^{N}K\binom{N}{K}p^{K}(1-p)^{N-K}=Np$. In general, the expected number of FTGs requiring retransmission in the $i$-th round of retransmission can be derived as $Np^{i}$. The likelihood of requiring the $i$-th round of retransmission corresponds to the probability that at least one FTG transmitted in that round experiences unrecoverable loss, which is given by $1-(1-p)^{Np^{i}}$. Using this, the expected total time for receiving all the data fragments across the required levels can be calculated as follows:
(2) $E[T_{total}]=t+\frac{nN-1}{r}+\sum\limits_{i=1}^{\infty}[1-(1-p)^{Np^{i-1}}](t+\frac{nNp^{i}-1}{r})$
Since we always have $10<p<$ 1, when $i\to\infty$, $[1-(1-p)^{Np^{i-1}}]\to 0$ and $(t+\frac{nNp^{i}-1}{r})\to t$. Consequently, the summation term in $E[T_{total}]$ diminishes, ensuring that $E[T_{total}]$ converges as $i\to\infty$.
Now the key is to estimate $p$, the probability that the data fragments of an FTG cannot be recovered from the losses. If each FTG has $n$ fragments and the transmission rate is $r$ fragments per second, sending $n$ fragments takes $n/r$ seconds. If the network latency for transmitting a single fragment from the sender to the receiver is $t$, the total time from when the first fragment of the FTG is sent to when the last fragment is received can be expressed as $T=t+(n-1)/r$. The total number of fragments sent during $T$ is
(3) $u=\frac{T}{1/r}=\frac{t+(n-1)/r}{1/r}=rt+n-1$
Assume the packet loss rate remains constant or exhibits minimal variation over short time intervals. Under this assumption, packet losses can be effectively modeled using a Poisson distribution. Specifically, the probability of exactly $j$ out of the $u$ fragments being lost within $T$ is given by:
(4) $Pr(v=j)=\frac{(\lambda T)^{j}e^{-\lambda T}}{j!}=\frac{{\lambda(t+\frac{n-1}{r})}^{j}e^{-\lambda(t+\frac{n-1}{r})}}{j!}$
These $j$ lost fragments may belong to different FTGs. If each FTG has $m$ parity fragments, given that there are $j$ ($j>m$) fragment losses during $T$, the probability that the lost data fragments in an FTG cannot be recovered is given by:
(5) $Pr(\text{unrecoverable}|v=j)=\frac{\sum\limits_{w=m+1}^{n}\binom{n}{w}\binom{u-n}{j-w}}{\binom{u}{j}}$
Therefore, the overall probability that the lost data fragments in an FTG cannot be recovered is given by:
(6) $\begin{split}p=&\sum\limits_{j=m+1}^{u}Pr(\text{unrecoverable}|v=j)Pr(v=j)\\ =&\sum\limits_{j=m+1}^{u}\frac{\sum\limits_{w=m+1}^{n}\binom{n}{w}\binom{u-n}{j-w}}{\binom{u}{j}}\frac{{\lambda(t+\frac{n-1}{r})}^{j}e^{-\lambda(t+\frac{n-1}{r})}}{j!}\\ =&\sum\limits_{j=m+1}^{rt+n-1}\frac{\sum\limits_{w=m+1}^{n}\binom{n}{w}\binom{rt-1}{j-w}}{\binom{rt+n-1}{j}}\frac{{\lambda(t+\frac{n-1}{r})}^{j}e^{-\lambda(t+\frac{n-1}{r})}}{j!}\end{split}$
However, when the packet loss rate is high (e.g., when each FTG experiences more than one fragment loss on average) and number of parity fragments per FTG is small (e.g., $m=0$ or $m=1$), unrecoverable data losses across groups become correlated, meaning multiple FTGs within $T$ are likely to experience unrecoverable losses simultaneously. This correlation invalidates the probability $p$ calculated using Equation 6, rendering it unsuitable for use in Equation 2 to compute $E[T_{total}]$. To address this issue, we adopt an alternative approach to calculate $p$ that accounts for the correlations introduced by high packet loss rates. This refined calculation ensures that Equation 2 remains a reliable approximation for the total transmission time, even under conditions of elevated packet loss. Given that $\frac{t+(n-1)/r}{n/r}$ FTGs are in transit during $T$, the average fragment losses each of these FTGs experiences is $\lambda T/(\frac{t+(n-1)/r}{n/r})=\frac{\lambda n}{r}$. An FTG experiences unrecoverable data loss only when more than $m$ fragments within it are lost. Therefore, the probability $p$ that an FTG experiences unrecoverable data losses can be calculated as follows:
(7) $\begin{split}p=&Pr(\text{more than}\ m\ \text{fragment losses in an FTG})\\ =&1-\sum\limits_{j=0}^{m}\frac{{(\frac{\lambda n}{r})}^{j}e^{-(\frac{\lambda n}{r})}}{j!}\end{split}$
In summary, the following optimization problem needs to be solved:
[…]
No matter which equation is used to calculate $p$, increasing $m$ has two opposing effects on $E[T_{total}]$: 1) As $N=\frac{S}{(n-m)s}$, a larger $m$ increases the number of FTGs, leading to more fragments transmitted during both the initial transfer and potential retransmissions, thus increasing $E[T_{total}]$. 2) A larger $m$ reduces the probability $p$ of unrecoverable data loss within an FTG, decreasing the likelihood of retransmissions and mitigating the increase in $E[T_{total}]$ caused by a higher $N$. As a result, an optimal value of $m$ exists that minimizes the expected total transmission time $E[T_{total}]$. In practice, since $p$ is relatively small, $E[T_{total}]$ usually converges for $i>50$. Given that $m$ is an integer and $m\leq n/2$ (as the number of parity fragments must not exceed the number of data fragments), finding the optimal $m$ that minimizes $E[T_{total}]$ is computationally straightforward. This can be achieved through an exhaustive search or a lightweight numerical optimization approach.
3.2.2. Minimizing Errors in the Received Data with a Guaranteed Data Transmission Time
In certain scenarios, scientists prioritize receiving data rapidly within a deterministic timeframe, even if it contains some level of error, rather than waiting for a prolonged and unpredictable period to obtain highly accurate data. For example, in scientific visualization, quickly rendering preliminary representations of large datasets allows researchers to identify trends, detect anomalies, or formulate hypotheses, even if the data is imperfect.
To ensure predictable transmission times, retransmissions must be minimized to avoid introducing non-deterministic delays. As a result, fault tolerance relies primarily on parity fragments generated by erasure coding. The key challenge is to determine which levels to transmit and the appropriate number of parity fragments created for the FTG at each level, so that the total transmission time stays within the user-specified timeframe while minimizing errors in the received data. The optimization model presented below is developed to address this challenge.
Assume the size of the $j$-th level of the refactored data is $S_{j}$, and $m_{j}$ parity fragments are generated for this level. Consequently, the number of FTGs required for the first level can be expressed as: $N_{j}=\frac{S_{j}}{(n-m_{j})s}$. If the first $l$ out of $L$ levels of the refactored data are sent without retransmission, the total transmission time can be calculated as:
(9) $\begin{split}T_{total}=&t+\frac{n(N_{1}+N_{2}+\ldots+N_{l})-1}{r}\\ =&t+\frac{n\sum\limits_{j=1}^{l}\frac{S_{j}}{(n-m_{j})s}-1}{r}\end{split}$
Let the user-specified constraint on total transmission time be $\tau$. Given that $20\leq m_{j}\leq n/$ 2, the range of $l$ can be roughly estimated by solving the following inequality:
(10) $t+\frac{n\sum\limits_{j=1}^{l}\frac{S_{j}}{(n-0)s}-1}{r}\leq T_{total}\leq t+\frac{n\sum\limits_{j=1}^{l}\frac{S_{j}}{(n-\frac{n}{2})s}-1}{r}\leq\tau$
To estimate errors in the received data, let $p_{1}$ denote the probability that an FTG in the first level experiences unrecoverable data loss. Assuming such losses are independent across groups, the probability that the first level cannot be recovered is $1-(1-p_{1})^{N_{1}}$. Since the first level is essential for data reconstruction, its loss results in a relative error of $\varepsilon_{0}=1$. If the first level is successfully recovered but the second level is not, the data can still be reconstructed with an error of $\varepsilon_{1}$, occurring with probability $(1-p_{1})^{N_{1}}[1-(1-p_{2})^{N_{2}}]$. Similarly, if the first two levels are recovered but the third fails, the error is $\varepsilon_{2}$, with probability $(1-p_{1})^{N_{1}}(1-p_{2})^{N_{2}}[1-(1-p_{3})^{N_{3}}]$. Following this pattern, for $l$ transmitted levels, the expected error of the reconstructed data is:
(11) $\begin{split}E[\varepsilon]=&[1-(1-p_{1})^{\frac{S_{1}}{(n-m_{1})s}}]\varepsilon_{0}+\\ &\sum\limits_{i=2}^{l-1}\prod\limits_{j=1}^{i-1}(1-p_{j})^{\frac{S_{j}}{(n-m_{j})s}}[1-(1-p_{i})^{\frac{S_{i}}{(n-m_{i})s}}]\varepsilon_{i-1}+\\ &\prod\limits_{j=1}^{l}(1-p_{j})^{\frac{S_{j}}{(n-m_{j})s}}\varepsilon_{l}\end{split}$
As previously discussed, the method for calculating $p_{j}$ depends on the packet loss rate. When the packet loss rate is relatively low, $p_{j}$ is calculated using Equation 6. Conversely, when the packet loss rate is high and correlations between FTGs become significant, $p_{j}$ is determined using Equation 7.
In summary, the complete optimization model, combining all components, is formulated as a nonlinear integer programming problem as follows:
(12) $\begin{split}\operatorname*{arg\,min}_{m_{j},\forall j\in\{1,\ldots,l\}}&\quad[1-(1-p_{1})^{\frac{S_{1}}{(n-m_{1})s}}]\varepsilon_{0}+\\ &\quad\sum\limits_{i=2}^{l-1}\prod\limits_{j=1}^{i-1}(1-p_{j})^{\frac{S_{j}}{(n-m_{j})s}}[1-(1-p_{i})^{\frac{S_{i}}{(n-m_{i})s}}]\varepsilon_{i-1}+\\ &\quad\prod\limits_{j=1}^{l}(1-p_{j})^{\frac{S_{j}}{(n-m_{j})s}}\varepsilon_{l}\\ \text{s.t.}&\quad t+\frac{n\sum\limits_{j=1}^{l}\frac{S_{j}}{(n-m_{j})s}-1}{r}\leq\tau\end{split}$
This model can be solved using numerical solvers such as SCIP (Bolusani et al., 2024) or other optimization tools capable of handling nonlinear integer programming.
4. Adaptive Data Transfer Protocols
In real-world networks, packet loss rates often fluctuate over time, introducing variability that can impact data transfer performance. Consequently, the fault-tolerance configurations derived from optimization models based on constant packet loss rates may become suboptimal, potentially degrading transmission efficiency or violating the user-specified constraints on total transmission time. To address this challenge, we integrate adaptive mechanisms into the data transfer protocols, which can dynamically adjust fault-tolerance configurations in response to changing packet loss rates, ensuring that the two user requirements outlined in Section 3.2 are consistently satisfied.
4.1. Data Transfer with Guaranteed Error Bound
The protocol for transferring data between sender and receiver with guaranteed error bound is shown in algorithm 1.
Sender
       Input: $\varepsilon$, $s$, $t$, $n$, $r_{ec}$, $r_{link}$, $\lambda$
       Parity Generation Thread
             Determine $l$ such that $\varepsilon_{l}\leq\varepsilon<\varepsilon_{l-1}$;
             Divide $l$ levels into $d=\frac{\sum_{i=1}^{l}S_{i}}{s}$ data fragments;
             Compute $r=min\{r_{ec},r_{link}\}$;
             Solve Equation 8 to obtain $m$;
             foreach $k$ ($k=n-m$) data fragments do
                   Generate $m$ parity fragments to form an FTG;
                   Store data and parity fragments of this FTG to $buf$;
                   if updated $\lambda$ notified by Receiver then
                         Update $m$ by solving Equation 8 again;
      Transmission Thread
             while $d>0$ do
                   Transmit fragments of an FTG in $buf$;
                   $d\leftarrow d-k$;
            Notify Receiver transmission ended;
             while True do
                   Wait for $lost\_FTGs\_list$ from Receiver;
                   if $lost\_FTGs\_list$ is empty then
                         break
                  foreach $lost\_FTG$ in $lost\_FTGs\_list$ do
                         Obtain fragments of $lost\_FTG$ from $buf$ and transmit;
                  Notify Receiver transmission ended;
Receiver
       Input: $T_{W}$
       Initialize $t_{start}\leftarrow t_{0}$, $lost\leftarrow 0$;
       while True do
             Receive fragments from Sender;
             foreach $FTG$ do
                   if fragment loss occurs then
                         Increment $lost$ by number of lost fragments;
                         if data fragment loss occurs then
                               if number of lost fragments $\leq m$ then
                                     Recover lost data fragments;
                              else
                                     Add $FTG$ to $lost\_FTGs\_list$;
            if $t_{now}-t_{start}\geq T_{W}$ then
                   Update $\lambda\leftarrow\frac{lost}{T_{W}}$ and notify Sender;
                   $t_{start}\leftarrow t_{now}$, $lost\leftarrow 0$
            if transmission ended notified by Sender then
                   Send $lost\_FTGs\_list$ to Sender;
                   if $lost\_FTGs\_list$ is empty then
                         break
Algorithm 1 Data Transfer with Guaranteed Error Bound
The sender is designed with multithreading capabilities to manage parity generation and data transmission simultaneously. In particular, the parity generation thread begins by analyzing the user-specified error bound to determine the appropriate number of levels for transmission. It then partitions the data into fragments and calculates the required number of parity fragments ($m$) needed for each FTG. This calculation involves solving the optimization problem defined in Equation 8, which depends on parameters such as the actual packet transmission rate ($r$) and the packet loss rate ($\lambda$) in lost packets per second. Notably, the actual packet transmission rate is determined as the minimum of the parity fragment generation rate ($r_{ec}$) and the network link transmission rate ($r_{link}$) for UDP packet of size $s$. The parity generation thread continuously generates $m$ parity fragments for every $k$ data fragments to form an FTG while monitoring for updates to $\lambda$ from the receiver. If the packet loss rate changes, the thread dynamically recalculates $m$, ensuring the system adapts to current network conditions. Simultaneously, the transmission thread handles the actual sending of fragments in FTGs. Upon completing the transmission of all data fragments, it notifies the receiver and waits for feedback on any lost FTGs. The thread then retransmits the lost FTGs iteratively until list of lost FTGs returned by the receiver is empty.
The receiver operates continuously to handle incoming fragments from the sender. For each FTG, it verifies if any fragment loss has occurred and counts the number of lost fragments within a user-specified time window, $T_{W}$. If data fragment loss is detected and the number of lost fragments in the FTG does not exceed $m$, the receiver recovers the missing data fragments using the remaining data and parity fragments. If recovery is not possible, the FTG is marked as unrecoverable and added to a list of lost FTGs. Once $T_{W}$ has elapsed, the receiver calculates the updated packet loss rate, $\lambda$ based on the total fragment losses within the time window and notifies the sender with the new $\lambda$. Additionally, when notified by the sender that the transmission has concluded, the receiver provides the sender with the list of lost FTGs for retransmission.
4.2. Data Transfer with Guaranteed Time
The protocol for transferring data between sender and receiver with guaranteed transmission time is shown in algorithm 2.
Sender
       Input: $\tau$, $[S_{1},S_{2},\ldots,S_{L}]$, $[\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{L}]$, $s$, $t$, $n$, $r_{ec}$, $r_{link}$, $\lambda$
       Parity Generation Thread
             Compute $r=min\{r_{ec},r_{link}\}$;
             Find all possible $l$ values by solving Equation 10;
             if no possible $l$ is found due to small $\tau$ then
                   Throw an exception;
            $\varepsilon_{min}\leftarrow 1.0$;
             foreach possible $l$ value $l^{\prime}$ do
                   Obtain $[m_{1},m_{2},\ldots,m_{l^{\prime}}]$ by solving Equation 12;
                   $\varepsilon\leftarrow$ value of Equation 11;
                   if $\varepsilon<\varepsilon_{min}$ then
                         $\varepsilon_{min}\leftarrow\varepsilon$;
                         $l\leftarrow l^{\prime}$ and $M\leftarrow[m_{1},m_{2},\ldots,m_{l^{\prime}}]$;
            Calculate the total number of data fragments $d=\frac{\sum_{i=1}^{l}S_{i}}{s}$;
             foreach $i\in\{1,2,\ldots,l\}$ do
                   Divide the $i$-th level into $d_{i}=\frac{S_{i}}{s}$ data fragments;
                   foreach $k_{i}$ ($k_{i}=n-m_{i}$) data fragments do
                         Generate $m_{i}$ parity fragments to form an FTG;
                         Store data and parity fragments of this FTG to $buf$;
                         if updated $\lambda$ notified by Receiver then
                               Update $\{m_{i},\ldots,m_{l}\}$ by solving Equation 12 again;
      Transmission Thread
             while $d>0$ do
                   Transmit fragments of an FTG in $buf$;
                   Extract $m_{i}$ from the fragments’ metadata;
                   $d\leftarrow d-(n-m_{i})$;
            Notify Receiver transmission ended;
Receiver
       Input: $T_{W}$
       Initialize $t_{start}\leftarrow t_{0}$, $lost\leftarrow 0$;
       while True do
             Receive fragments from Sender;
             foreach $FTG$ do
                   Extract $m_{i}$ from the fragments’ metadata;
                   if fragment loss occurs then
                         Increment $lost$ by number of lost fragments;
                         if data fragment loss occurs then
                               if number of lost fragments $\leq m_{i}$ then
                                     Recover lost data fragments;
                              else
                                     Mark the $i$-th level as unrecoverable;
            if $t_{now}-t_{start}\geq T_{W}$ then
                   Update $\lambda\leftarrow\frac{lost}{T_{W}}$ and notify Sender;
                   $t_{start}\leftarrow t_{now}$, $lost\leftarrow 0$
            if transmission ended notified by Sender then
                   break
Algorithm 2 Data Transfer with Guaranteed Time
Similar to algorithm 1, the sender in algorithm 2 leverages multithreading to manage parity generation and data transmission concurrently. The parity generation thread begins by calculating the effective packet transmission rate ($r$), determined as the minimum of the parity fragment generation rate ($r_{ec}$) and the network link transmission rate ($r_{link}$). Using this transmission rate, it solves Equation 10 to identify all feasible values for number of levels that can be transmitted within the given deadline $\tau$. If no feasible value exists, the protocol raises an exception, indicating that the specified deadline is too stringent for the current system configuration. Once all feasible values are identified, the parity generation thread refines the selection by determining the optimal number of levels $l$ to transmit and calculating the most effective combination of redundancy parameters $[m_{1},m_{2},\ldots,m_{l}]$ for each level. This is achieved by solving the optimization problem defined in Equation 12, which minimizes the overall error bound ($\varepsilon$) while satisfying the timing and transmission constraints. Each level is then divided into data fragments of size $s$. For the $i$-th level, the parity generation thread generates $m_{i}$ parity fragments for every $k_{i}$ ($k_{i}=n-m_{i}$) data fragments to form an FTG. The parity generation thread is designed to dynamically adapt to changes in network conditions. If the receiver notifies the sender of an updated packet loss rate ($\lambda$), the thread recalculates the redundancy parameters $m_{i}$ by solving Equation 12. However, FTGs that have already been transmitted are excluded from consideration, and the time elapsed since the transmission began is deducted from the remaining deadline $\tau$ when solving Equation 12.
The receiver continuously processes incoming fragments and monitors fragment losses. For each FTG, it extracts the level-specific redundancy parameter ($m_{i}$) from the fragments’ metadata to determine recoverability. If data fragments are lost and the total number of lost fragments does not exceed $m_{i}$, the receiver restores the missing fragments. Otherwise, the affected level is marked as unrecoverable. The receiver also maintains a time window ($T_{W}$) to calculate the packet loss rate ($\lambda$) and periodically notifies the sender of updated $\lambda$ values. Once the sender signals the end of transmission, the receiver concludes its operations.
5. Evaluation
Simulation serves as the primary evaluation method for the proposed adaptive data transfer protocols due to its ability to provide a controlled and flexible environment. Unlike real networks, where packet loss rates are unpredictable and difficult to control, simulation allows systematic exploration of protocol behavior under diverse and extreme packet loss conditions. To complement the simulation-based evaluation, we also conducted real network testing focused on validating the functionality of the proposed protocols in practical scenarios.
5.1. Testing Data and Network
The scientific data used in our evaluation was generated by the Nyx cosmology simulation (NYX simulation, 2019) and refactored into a hierarchical representation with four levels using the pMGARD (Liang et al., 2021) library. The sizes of the four levels were $S_{1}=668$ MB, $S_{2}=2.67$ GB, $S_{3}=5.42$ GB, and $S_{4}=17.99$ GB. The relative L-infinity errors for reconstructing the data using the first $i$ levels of the hierarchy, denoted as $\varepsilon_{i}$, were set as follows: $\varepsilon_{1}=0.004$, $\varepsilon_{2}=0.0005$, $\varepsilon_{3}=0.00006$, and $\varepsilon_{4}=0.0000001$. Note that we do not evaluate data refactoring in this work, as it is typically performed during the data generation process and stored on the system where it is produced before any transfer occurs. Evaluations of data refactoring techniques can be found in existing works, such as (Liang et al., 2021; Wan et al., 2023). The testing network was configured with the sender deployed on a workstation at our university and the receiver operating on a virtual machine hosted by CloudLab (CloudLab, 2025), a cloud infrastructure that is part of the National Science Foundation’s NSFCloud program.
5.2. Simulation-Based Evaluation
5.2.1. Design of the Simulation Model
We developed a discrete-event simulation framework using SimPy (SimPy, 2025), a Python library for modeling discrete-event systems. The core simulation models key operations of data transfer protocols. A process models the sender, which continuously generates and transmits packets. For every $k$ data fragments, the sender generates $m$ parity fragments, transmitting each packet every $1/r$ seconds, where $r$ is the transmission rate. Each packet experiences a transmission latency of $t$ seconds. If the packet is not marked as lost by the packet loss process, it is added to a FIFO queue. Another process models the receiver, which retrieves packets from the FIFO queue. The packet loss process simulates losses by generating random time intervals between losses. When a loss event occurs, the packet is marked as lost if the loss event queue is not empty. Afterward, the loss event queue is cleared. A separate process handles control message exchanges between the sender and receiver.
Our framework also extends the core simulation to model specific protocols. For TCP, parity fragment generation is disabled, and acknowledgment and retransmission mechanisms are simulated. For UDP-based protocols with erasure coding, the receiver analyzes metadata to identify FTGs with data losses and checks if recovery is possible. If the required data levels are insufficient and an error bound must be guaranteed, the receiver requests retransmissions until all necessary fragments are received. For the adaptive data transfer protocols, the receiver periodically measures the packet loss rate and updates the sender, which then recalculates the optimal $m$ values and adjusts parity generation to ensure efficient, reliable data transfer.
5.2.2. Simulation Parameter Settings
To determine the simulation parameters, we conducted extensive measurements of data transmission over the testing network. First, we measured the transmission latency $t$ by sending 4,096-byte UDP data fragments for 15 minutes and recording the latency of each fragment. These tests were repeated across multiple days to account for temporal variations. The results showed stable UDP transmission latency ranging from 0.0095 to 0.015 seconds, so we set $t=0.01$ seconds for our simulation.
Next, we measured the maximum network transmission rate $r_{link}$ and the parity fragment generation rate $r_{ec}$. Using 4,096-byte packets, we found that $r_{link}=19,144$ packets per second. To determine $r_{ec}$, we used the “liberasurecode” library (Unified Erasure Coding interface for common storage workloads, 2022) to generate parity fragments for FTGs, each consisting of $n=32$ fragments. As $m$ (the number of parity fragments) increased from 1 to 16, $r_{ec}$ decreased from 319,531 to 41,561 fragments per second. Since $r_{ec}$ remained higher than $r_{link}$ even at $m=16$, the actual transmission rate was limited by the network link, $r=r_{link}$. However, in high-bandwidth networks where $r_{link}>r_{ec}$, transmission could be limited by $r_{ec}$.
We determined packet loss rates based on common wide-area network classifications: low, medium, and high. A low loss rate (under 0.1%) corresponds to $\lambda_{low}=r\times 0.1\%=19$ lost packets per second. A medium loss rate (2%) results in $\lambda_{medium}=r\times 2\%=383$ lost packets per second, while a high loss rate (5%) gives $\lambda_{high}=r\times 5\%=957$ lost packets per second. The packet loss generation process uses an exponential distribution with these rate parameters to model the time intervals between consecutive packet loss events.
To model time-varying packet loss rates, we use a hidden Markov model (HMM), as suggested in studies such as (Rodbro et al., 2006; Rossi et al., 2006; Silveira and de Souza e Silva, 2007). The HMM consists of three states—low, medium, and high—each associated with a Gaussian distribution. Specifically, the low state has $\mu=19$, $\sigma=2$, the medium state has $\mu=383$, $\sigma=40$, and the high state has $\mu=957$, $\sigma=100$. Transitions between states follow a continuous-time Markov chain with an exponential holding time distribution, where the rate parameter is 0.04, meaning state transitions occur approximately every 25 seconds. Within each state, the packet loss rate is randomly generated based on the corresponding Gaussian distribution, simulating more realistic, time-varying network conditions.
Finally, protocol-specific parameters are configured using commonly accepted values. For TCP, the retransmission timeout is set to twice the transmission latency, and the duplicate ACK threshold for triggering fast retransmission is set to 3. For the adaptive data transfer protocols, the time window for measuring and updating the packet loss rate is set to 3 seconds, balancing responsiveness to changing network conditions with measurement stability.
5.2.3. Data Transfer under Static Packet Loss Rates
(a) Low packet loss rate
(b) Medium packet loss rate
(c) High packet loss rate
Figure 2. Total time for transferring data with guaranteed error bound under static packet loss rates
While real-world network packet loss rates fluctuate over time, we conducted simulation-based evaluations under static packet loss rates to validate the correctness of the models we developed in Section 3.
First, we evaluated the data transfer process for ensuring an error-bound guarantee on the received data. Assuming the user specifies that the error in the received data must remain below $0.00001$, all four levels of the refactored NYX cosmology simulation data must be successfully received to meet this requirement. We simulated the transfer of these four levels using the TCP protocol. We also simulated an alternative UDP-based protocol that incorporates erasure coding with a static number of parity fragments for each FTG, denoted as $m$, alongside passive retransmission to ensure fault tolerance. For each possible value of $m$, we conducted simulations, recording the total time required to complete the transfer of all four levels. Furthermore, we computed the expected total transfer time, $E[T_{total}]$, using the equations presented in Section 3 and compared the calculated results with the simulation outcomes. The evaluation results are presented in Fig. 2, providing the following insights: 1) When the TCP protocol is used, the data transmission time increases significantly as the packet loss rate rises. 2) For the UDP-based protocol that integrates erasure coding and passive retransmission, the total transmission time derived from the theoretical equations in Section 3 aligns closely with the simulation results. This alignment confirms that our model effectively captures the impact of packet loss, parity fragment generation, and retransmission on overall transmission performance. 3) When the packet loss rate is low (Fig. 2(a)), increasing the number of parity fragments per FTG negatively affects performance due to the added overhead. However, as the packet loss rate increases (Fig. 2(b) and Fig. 2(c)), there exists an optimal value of $m$ that minimizes the total transmission time.
(a) Low packet loss rate
(b) Medium packet loss rate
(c) High packet loss rate
Figure 3. Error bounds of data received within guaranteed transmission time under static packet loss rates
Second, we evaluated the data transfer process under a guaranteed transmission time constraint, which enables users to receive data within a specified time frame while accepting the trade-off that the error bound of the received data may not be strictly guaranteed. As shown in Fig. 2, the minimum transmission times required to transfer all levels of the refactored data under different packet loss rates were 378.03 seconds for $\lambda=19$, 401.11 seconds for $\lambda=383$, and 429.75 seconds for $\lambda=957$. These values were subsequently used as the transmission time constraint $\tau$ in solving the optimization model presented in Equation 12. The optimization results indicate that to minimize the error bound of the received data, the optimal number of parity fragments per FTG at all four levels should be set as follows: for $\lambda=19$, the values are $m_{1}=5$, $m_{2}=4$, $m_{3}=2$, and $m_{4}=0$; for $\lambda=383$, the values are $m_{1}=8$, $m_{2}=7$, $m_{3}=7$, and $m_{4}=0$; and for $\lambda=957$, the values are $m_{1}=12$, $m_{2}=11$, $m_{3}=11$, and $m_{4}=0$. To evaluate the effectiveness of these fault-tolerance configurations, we conducted simulations using the derived configurations, running each scenario 100 times to account for the nondeterministic nature of data errors. We recorded the number of instances where the received data satisfied different error bounds, such as $\varepsilon_{0}$, $\varepsilon_{1}$, and so on. This approach allowed us to quantify the frequency of achieving varying levels of data accuracy under the given configurations. Additionally, to highlight the advantages of the optimized configurations, we performed simulations with alternative configurations where the number of parity fragments per FTG was set uniformly across all four levels. As illustrated in Fig. 3, the optimized configurations consistently ensured that the transmission time remained within the specified constraint while achieving a high likelihood of maintaining a low error bound ($\varepsilon_{3}$) in the received data. In contrast, the alternative configurations either met the time constraint but failed to recover the data entirely (resulting in an error bound of $\varepsilon_{0}$), or exceeded the specified transmission time constraint.
5.2.4. Data Transfer under Time-Varying Packet Loss Rates
When the packet loss rate changes over time, a static fault-tolerance configuration fails to achieve optimal transmission performance. We conducted simulation-based evaluations to assess the effectiveness of the proposed data transfer protocols in dynamically adjusting fault-tolerance configurations in response to time-varying packet loss rates.
First, we conducted simulations to evaluate the performance of three data transfer approaches under a guaranteed error bound: the TCP protocol, a UDP-based protocol incorporating erasure coding with a static number of parity fragments per FTG, and the proposed adaptive protocol shown in algorithm 1. As illustrated in Fig. 4, the adaptive protocol consistently outperforms the other two approaches under time-varying packet loss rates. Notably, the adaptive protocol completes the data transmission in 388.8 seconds, reducing the transmission time by 30 seconds compared to the minimum time achieved by the static fault-tolerance configuration.
Figure 4. Total time for transferring data with guaranteed error bound under time-varying packet loss rates
Second, we conducted simulations to evaluate the proposed protocol in algorithm 2 for transferring data within a guaranteed time constraint under time-varying packet loss rates. Since the minimum transmission time for all data levels using the protocol in algorithm 1 under time-varying packet loss rates is 388.8 seconds, we selected this duration as the time constraint. This choice reflects a scenario where the user is willing to trade some accuracy of the received data for potentially reduced transmission time. To determine the optimal fault-tolerance configuration under this constraint, we solved the optimization model in Equation 12 using static packet loss rates. The results indicate that to meet the time constraint while minimizing the error bound, the optimal number of parity fragments per FTG at all four levels should be configured as follows: for $\lambda=19$, the values are $m_{1}=9$, $m_{2}=6$, $m_{3}=4$, and $m_{4}=0$; for $\lambda=383$, the values are $m_{1}=16$, $m_{2}=8$, $m_{3}=0$, and $m_{4}=0$; and for $\lambda=957$, the values are $m_{1}=15$, $m_{2}=9$, $m_{3}=0$, and $m_{4}=0$. We performed 100 simulation runs using these static fault-tolerance configurations under time-varying packet loss rates. Additionally, we ran another 100 simulations using the adaptive protocol in algorithm 2 under the same conditions, recording the number of instances where the received data met various error bounds. As shown in Fig. 5, both the static and adaptive fault-tolerance configurations successfully met the time constraint. However, the adaptive configuration consistently resulted in received data satisfying lower error bounds more frequently than the static configurations, demonstrating its effectiveness in improving data accuracy while maintaining the required transmission time under fluctuating packet loss rates.
Figure 5. Error bounds of data received within guaranteed transmission time under time-varying packet loss rates
5.3. Real Network Testing
5.3.1. C++ Implementation of Adaptive Data Transfer Protocols
To evaluate the performance of data transmission over real wide-area networks, we developed a prototype of the proposed adaptive data transfer protocols, as described in algorithm 1 and algorithm 2, using C++. Our implementation leverages the Boost.Asio (Boost.Asio, 2025) library for efficient UDP-based data transmission. Additionally, we utilize the Protocol Buffers (Protobuf) library (Protocol Buffers, 2025) to structure packet contents, embedding essential erasure coding metadata—such as the level and FTG each fragment belongs to, as well as its corresponding redundancy level ($m$)—alongside the actual data or parity fragment. The prototype, comprising over 1,600 lines of C++ code, provides a scalable and flexible foundation for evaluating the proposed adaptive transfer mechanisms in real-network environments.
5.3.2. Testing Results
Figure 6. Total time for transferring data with guaranteed error bound using different data transfer protocols
In each test run, we conducted several measurements. First, we used the native TCP protocol to transfer all levels of the refactored Nyx data and recorded the total transmission time. Next, we employed the Globus (Globus, 2025) data transfer service to transfer the same data and recorded the transmission time. Then we used the adaptive protocol described in algorithm 1 to transmit the data, ensuring an error bound guarantee of $\varepsilon_{4}$ and requiring the successful receipt of all levels. Finally, we applied the adaptive protocol in algorithm 2 to transfer the data within a fixed time constraint, set to 90% of the transmission time achieved by the protocol in algorithm 1. To account for variations in packet loss rates, we conducted five test runs at different times and on different days.
Test ID Time Constraint(s) Error Bound
1 367.06 $\varepsilon_{2}$
2 392.08 $\varepsilon_{2}$
3 339.32 $\varepsilon_{2}$
4 371.08 $\varepsilon_{2}$
5 348.66 $\varepsilon_{1}$
Table 2. Error bounds of data received within guaranteed transmission time
As shown in Fig. 6, when all levels of the Nyx data must be received, the transmission time for native TCP and Globus varies significantly across the five test runs, emphasizing their sensitivity to packet loss. In contrast, our adaptive data transfer protocol not only reduces transmission time substantially but also offers greater stability. Furthermore, as demonstrated in Table 2, when using the adaptive protocol described in algorithm 2 to meet the selected time constraints, four out of five test runs achieved data within the error bound $\varepsilon_{2}$, while one run resulted in data within the error bound $\varepsilon_{1}$. This shows that by utilizing algorithm 2, we can guarantee a shorter transmission time, albeit with a slight sacrifice in the accuracy of the received data.
6. Related Work
UDP, a connectionless and low-latency transport protocol, lacks built-in error correction mechanisms (Postel, 1980), making it particularly susceptible to packet loss in high-performance and wide-area network environments. To address this limitation, many studies have explored lightweight and adaptive flow control mechanisms to improve the reliability of UDP-based data transfer. One notable approach is Performance-Adaptive UDP (PA-UDP), a high-speed data transfer protocol proposed by (Eckart et al., 2008). PA-UDP enables dynamic flow control based on the sending system’s performance characteristics, allowing UDP packets to be sent asynchronously without blocking TCP transmission. The SDUDP framework (Wang et al., 2017) takes advantage of Software-Defined Networking (SDN)’s capabilities, such as centralized flow control and real-time packet monitoring, to achieve reliable UDP data transfer. Beyond SDUDP, various studies have explored Reliable UDP (RUDP) protocols designed to enhance UDP’s performance while maintaining its low-latency advantage over TCP. These protocols typically introduce mechanisms such as selective acknowledgments, forward error correction (FEC), and retransmission strategies to mitigate packet loss. However, many RUDP implementations require modifications to end-host systems, making deployment challenging in large-scale networks (Tran and Choi, 2007; Masirap et al., 2016; Wang et al., 2006). In addition, QUIC (Quick UDP Internet Connections), a transport layer protocol developed by Google, uses mechanisms like congestion control and loss recovery to enhance packet delivery (Langley et al., 2017).
To further optimize UDP-based data transfer, research has focused on detecting and predicting packet loss more efficiently. For instance, Wu et al. (Wu et al., 2022) proposed a novel framework, LossDetection, for real-time packet loss detection in both TCP and UDP traffic. LossDetection employs Feature-Sketch, a technique that extracts key features from packet streams for efficient and scalable loss identification. By detecting packet loss patterns dynamically, this method enables more effective retransmission strategies. Another approach to mitigating packet loss involves predictive modeling. Giannakou et al. (Giannakou et al., 2020) developed a lightweight Random Forest Regression (RFR) model to predict retransmissions in scientific data transfers. Their model builds upon existing loss detection techniques by forecasting packet loss trends and proactively adjusting transmission parameters to minimize unnecessary retransmissions.
Beyond optimizing flow control mechanisms, another type of approach to improving UDP reliability involves introducing redundancy into transmitted data to tolerate packet losses. A fundamental technique used in these approaches is erasure coding, which has been extensively studied and implemented to enhance fault tolerance in both data storage and transmission. Foundational work by Reed and Solomon introduced Reed-Solomon codes (Wicker and Bhargava, 1999), which remain a gold standard for error correction. Several studies have explored the integration of erasure coding with UDP-based data transfer to improve reliability. For instance, Byers et al. (Byers et al., 1998) demonstrated that incorporating Reed-Solomon codes into UDP streaming significantly enhances packet recovery, mitigating losses common in lossy network conditions. Further research has investigated the use of fountain codes, such as Raptor codes, over UDP to facilitate reliable high-bandwidth data transfers. Mtibaa and Bestavros proposed RC-UDP (Mtibaa et al., 2018), a framework that leverages Raptor coding over UDP to achieve efficient and reliable data transmission. Their findings indicate that this approach effectively addresses packet loss without incurring significant overhead.
Despite these advancements, several challenges remain unaddressed. One major challenge is the computational and bandwidth overhead introduced by fault-tolerance techniques such as erasure coding. This overhead can be particularly significant when transferring large-scale scientific datasets over wide-area networks. To address this, we propose integrating erasure coding with lossy compression, effectively reducing the data volume while maintaining a controllable level of fidelity. Another key challenge is the dynamic nature of network conditions, particularly fluctuating packet loss rates, which can lead to suboptimal performance when using static fault-tolerance configurations. Existing static approaches fail to adapt to real-time variations in network reliability, often resulting in either excessive redundancy or insufficient protection against data loss. To overcome this, our work introduces an adaptive optimization framework that dynamically adjusts erasure coding parameters based on real-time network conditions.
7. Conclusion
In this paper, we introduced Janus, a resilient and adaptive data transmission approach designed to enhance the efficiency and reliability of scientific data transfers over wide-area networks. By leveraging UDP, integrating erasure coding for fault tolerance, and incorporating error-bounded lossy compression to reduce overhead, Janus enables users to optimize data transmission time while preserving accuracy. Furthermore, its ability to dynamically adjust erasure coding parameters based on real-time network conditions ensures sustained performance under fluctuating environments. Our extensive evaluations, including both simulations and real-network experiments, demonstrate that Janus significantly improves transfer efficiency while maintaining data integrity. As cross-facility scientific workflows continue to expand, solutions like Janus will play a crucial role in enabling timely and scalable data sharing, ultimately accelerating scientific discovery.